OpenStudy (anonymous):

Find the volume bounded by the cylinder x^2+y^2=25 and the planes y = z, x = 0, z = 0 in the first octant

OpenStudy (irishboy123):

you're looking at \[\iint f(x,y) \; dA = \iint y \; dx \; dy\] specifically \[\int\limits_{x= 0}^{5} \int\limits_{y=0}^{\sqrt{25 - x^2}} y \; dx \; dy\]